# Machine learning classification of black holes in the mass–spin diagram

### v2

@@ I Introduction
Figure 1 displays an HR diagram from a synthetic stellar population generated with the COMPAS rapid isolated-binary evolution code with default parameter settings [6]. The stellar classifications, as the HR diagram reveals, are intimately tied to the phases of stellar evolution through the stars’ lifespans. We filter this prototypical theoretical population to obtain a small randomized subset whose evolutionary stages are shown in Fig. 1. In chronological order, these intermediate-to-high mass stars begin core hydrogen fusion, i.e. the main sequence (red dashed lines) and move through a very brief Hertzsprung gap characterized by shell helium burning, which is followed by drastic radial expansion as core helium fusion ignites and constitutes the dominant component of the stellar luminosity in the giant branch (dot-dash purple lines). Then these stars lose their envelopes, either through binary mass transfer processes or stellar wind mass loss, revealing the core-helium burning phase of stellar evolution (dotted green lines) which leads imminently to gravitational collapse, possibly associated with a supernova explosion and compact object formation. A few surviving white dwarfs (gray dots) and neutron stars (blue stars) occupy their respective lower luminosity regions with the neutron stars effective temperature being much higher due to differences in their supernovae mechanisms. This highlights the utility of the HR diagram for organizing the variety of stellar types and diagnosing stellar evolution models. Extensions of the HR diagram are still being found for varied purposes [7, 8, 9, 3, 10], underscoring its impact in modern astronomy. However, black holes could only be drawn on the HR diagram if they are embedded in an accretion disk or due to non-classical possibilities such as Hawking radiation.
@@
Motivated by this paradigm, we ask: can an analogous diagram be constructed for astrophysical black holes? We propose that, similar to the HR diagram, one can visually explore how black holes can be organized and classified. While black holes are historically challenging to observe, their known macroscopic properties of mass and spin from General Relativity offer natural axes for classification. With electromagnetic (EM) techniques such as X-ray spectroscopy and radio interferometry providing constraints on stellar-mass and supermassive black holes [11, 12, 13, 14] and with gravitational-wave (GW) detections now delivering catalogs of stellar-mass black hole mergers [15, 16, 17], we can consider how patterns in black hole properties may emerge.
@@
Recent years have seen substantial effort in quantifying and interpreting the mass–spin relations of BHs across the mass spectrum. For stellar-mass black holes in X-ray binaries, two primary techniques—continuum-fitting and iron-line reflection spectroscopy—have yielded a range of spin magnitudes, underscoring the complexity of natal spin and accretion histories [18, 19]. At the supermassive end, X-ray reflection analyses show that the majority with masses below $\approx$3\text{\times}{10}^{7}$\rm\,M_{\odot}$ are rapidly spinning ($\chi\gtrsim$ 0.9), while those with higher masses exhibit a tentative trend toward lower spins [13]. The elusive intermediate-mass black holes ($\approx 10^{2}-10^{5}\rm\,M_{\odot}$) remain less certain, but numerous strong candidates are found in globular clusters, hyper-luminous X-ray sources, and dwarf-galaxy nuclei, albeit with large uncertainties [20, 21], and a few have been confidently detected via GWs, such as GW190521 [22] and GW231123 [23], calling into question the mass boundaries that typify this mass regime.
@@
Emergent gravitational-wave population studies suggest a mass-dependent spin distribution in merging binaries [24], and prior studies have explored black hole mass, spin, and redshift relationships in GW data [25, 26, 27]. An observed anti-correlation in the $\chi_{\rm eff}$$q$ plane [28, 29, 30], where $\chi_{\rm eff}$ is the aligned effective spin parameter [31, 32] and $q$ the mass ratio, was subsequently confirmed in more recent analyses of gravitational-wave catalogs and helps reveal subpopulations in the observed dataset associated with different formation scenarios [33, 34, 35]. Structures can be inferred from the spin distributions using data-driven models that capture unanticipated features without inserting systematic bias [36]. Investigations into the influence of stellar rotation on black hole mass–spin outcomes has long been studied and is model dependent [37, 38], which often involves implicitly projecting onto the mass–spin plane. These works significantly advance understanding of stellar-mass black hole populations, and motivate development of a tool for interpreting black hole mass and spin measurements.
@@
To begin sketching this framework, we present a cartoon schematic of the mass–spin plane alongside a physically motivated trajectory labeled “Cosmic Accretion,” representing a black hole analogue of the stellar main sequence. While there is only one main sequence for stars, we propose that black holes may have many depending on if and how they grow in mass and spin. We quantify this by sampling from physically motivated spin and mass distributions across the stellar-mass, intermediate-mass, and supermassive black hole regimes, and modeling their redshift evolution under Eddington-limited accretion to trace evolutionary tracks/trajectories across the diagram. Next, we construct a mass–spin diagram focused on the isolated-binary formation channel [39] of stellar-mass black holes using three state-of-the-art prescriptions for obtaining high black hole spin magnitude from tidal synchronization of its Wolf-Rayet progenitor’s spin, demonstrating the model-specific regions that a population of black holes can occupy in a canonical formation scenario of stellar-mass black holes.
@@
This paper is organized as follows. We present the theoretical motivation for the black hole mass-spin diagram and a couple examples to illustrate its utility in Sec. II. In Sec. III we apply machine learning classification pipelines to mass–spin datasets. In Sec. IV we summarize our analysis and findings, and in Sec. V we discuss the implications of our results with emphasis on future investigations.
Figure 2: A sketch of the mass–spin diagram for astrophysical black holes with dimensionless spin angular momentum (i.e. $\chi=c|\mathbf{S}|/Gm^{2}$) on the vertical axis. Regions are colored, as indicated in the legend, according to a selection of formation channels of black holes in single, binary, or higher systems across the mass spectrum: binary stars in isolation, dense stellar clusters, canonical population III stars, and hierarchical galaxy mergers. In reality, these formation channels will have overlap and their prevalence will depend on redshift. The two green vertical patches with red hatch symbolize the commonly used boundaries between the stellar-mass, intermediate-mass, and supermassive regimes. Typical astrophysical processes that drive the mass and spin growth across the diagram are written in each regime, with small arrows indicating uncertainty in the region borders for mass and spin predictions of formation channels. The solid blue line extending from the stellar-mass regime to the upper right corner of the diagram represents the direction of a Cosmic Accretion main sequence.
Figure 3: Theoretical evolutionary tracks of accreting black holes across the mass–spin diagram. Black holes are drawn from initial mass functions and simple natal spin prescriptions, and are evolved from redshift $z=10$ to $z=0$ along the trajectories. Stellar (blue dots and lines) and supermassive (red triangles and lines) black holes undergo an approximate thin disk evolution, while intermediate-mass black holes (green squares and lines) experience chaotic accretion modeled with a Gaussian drift. The flow of points toward the upper right corner of the panel is a quantitative example of a Cosmic Accretion main sequence (i.e., the blue solid line of Fig. 2).
@@ II Black hole mass-spin diagrams
Black holes cannot meaningfully be placed on an HR diagram as they lack intrinsic luminosity but a mass–spin diagram provides an analogous classification tool. In this space, every black hole occupies a point whose coordinates can shift over cosmic time due to formation, accretion, merger events, or even Hawking evaporation, thereby tracing out trajectories (or tracks) that reflect astrophysical and cosmological history. The mappings between redshift and environment dependent accretion states, and between pre- and post-merger binary component masses and spins implies the existence of black hole main sequence analogues for dominant growth pathways of mass and spin. Understanding such mappings allows inference of black hole evolutionary typology across cosmic time.
@@
Modern EM techniques have placed increasingly precise limits on spin magnitudes, both in X-ray binaries and quasars/AGN [12, 13, 58]. Measurements of X-ray binaries can in principle constrain the spin orientation of the accreting black hole but the orbital inclination introduces burdensome systematics. Further, it is an active area of research to develop new methods for EM observations to confidently identify dual- or multi-AGNs [59] and to distinguish a single supermassive black hole from a supermassive black-hole binary within a galactic nucleus [14], indicating a bright future for EM measurements of black hole spin orientations relative to a companion or other environmental aspects.
Black holes are observed confidently in two main mass regimes: stellar-mass black holes through electromagnetic methods, such as accreting binaries [12, 60] or astrometry, and through the terrestrial gravitational-wave detector network LVK [34]; and supermassive black holes in galactic nuclei and quasars, observed via multi-wavelength spectroscopy, gas and stellar dynamics, and reverberation mapping [61, 62, 13, 63, 64, 60], and hopefully via gravitational waves with a future space-based detector such as LISA via binary or minor mergers [65]. The intermediate-mass regime is populated with many strong candidates [20, 21] but has not yet been confidently confirmed beyond the few hundred solar mass extraordinary events from LVK. While these LVK events confirm the existence of black holes with masses greater than 100 $\rm\,M_{\odot}$ [66], it remains to be seen how the rest of the gap, ie from $\sim$1000 to $\sim 10^{5}$, is filled.
@@
Across the black-hole mass spectrum, several formation channels populate regions or trajectories in the mass–spin plane, which can be distinct or highly overlapping depending on the underlying assumptions. Stellar core collapse can produce stellar-mass black holes with low to moderate natal spins depending on the nature of angular momentum transport and possible tidal evolution during progenitor evolution [71, 72]. Binary mergers in dense environments lead to spin-up and mass growth via well-characterized remnant formulae. Accretion-driven evolution from both disc feeding and tidal disruptions can also significantly affect black hole spins depending on the Eddington fraction of the accretor and the tidal radius of the disruptor, respectively. Supermassive black holes grow through galaxy mergers and prolonged accretion episodes, resulting in mass–spin evolution that is skewed toward high spin in low-mass SMBHs but possibly limiting at high mass [73, 74, 75]. Also primordial black holes which, if they exist, could span a wide range of masses and spins and serve as fossils for cosmology.
@@
In Figure 2 we present a schematic mass–spin diagram summarizing some main theoretical possibilities and expected trajectories. Each shaded region corresponds to a specific black hole formation channel, such as isolated stellar binaries, stellar clusters, hierarchical galaxy mergers, or Population III stars, and spans a range of spin magnitudes and black hole masses predicted by each scenario. Horizontal and vertical transitions are overlaid to indicate astrophysical processes relevant for growing the black hole masses and/or spins. The processes for intermediate-mass black holes, which may be populated by multiple overlapping formation pathways, are expected to be a mixture of processes identified from the stellar (i.e. stellar collapse, spin inheritance, tidal spin-up of the progenitor, and accretion) and supermassive (dynamical friction, binary loss-cone stellar scattering, disk-driven migration and accretion) regimes [20, 21]. The commonly used boundaries between the three mass regimes are shown for reference in green regions with red hatching. The small arrows colored by formation channel indicate uncertainties on the theoretical boundaries of the formation channel. Therefore, this diagram serves as a visual analog to the HR diagram for stars: a two-dimensional landscape in which different populations occupy distinct, physically motivated regions, and evolutionary tracks and snapshots can be mapped and classified.
@@
We quantitatively illustrate this in Figure 3 which shows the redshift-dependent trajectories of black holes across the mass–spin plane under different physical growth scenarios, representing stellar-mass black holes (blue circles), intermediate-mass black holes (green squares), and supermassive black holes (red triangles). Each line, i.e. trajectory, traces the evolution of an individual black hole from redshift $z=10$ to $z=0$, with markers indicating the initial position in mass–spin space. The stellar-mass black holes are born with a power-law mass distribution ($m\sim m^{-2.3}$ [78]) and uniform spin magnitudes [79, 80], and evolve under an approximation to the classic model of [81] where we assume $z_{i}\rightarrow z_{j}$ with $i<j$ defines a redshift timestep. In this approximate model, a black hole of mass $m$ accretes an amount of mass $\Delta m(z)=m(z_{j})-m(z_{i})=m_{\rm n}(1+1.5(1-\bar{z}))$ where $\bar{z}=z/10$ and $m_{\rm n}$ is the natal mass, and undergoes corresponding spin increase via $\Delta\chi(z)=\chi_{i}+(\chi_{\rm eq}-\chi_{i})\eta\Delta m(z)/m$ where $\chi_{\rm eq}=0.998$ and $\eta=0.3$ is an ad-hoc parameter that controls the efficiency of spin-up. Intermediate-mass black holes are drawn from a log-flat mass distribution and bimodal initial spin distribution [82], and undergo modest stochastic mass growth $\Delta m(z)=m_{\rm n}(1+1.5(1-\bar{z}))(1+\mathcal{N}(0,0.003))$ accompanied with conditional spin drift (i.e., aggressive upward spin drift if $\chi<0.6$ at a given $z$), loosely modeling the combined effect of hierarchical mergers and chaotic accretion [83, 20]. We assume supermassive black holes originate from a 60%/40% mixture of light and heavy seed populations (light seeds from a Schechter-like distribution with natal spins $\sim\mathcal{N}(0.5,0.2)$ and heavy seeds from a log-uniform distribution with natal spins from $\sim\mathcal{N}(0.8,0.1)$) [84, 85, 74, 86, 87, 88, 89], and grow through coherent thin-disk accretion via $\Delta m(z)=m_{\rm n}(1+3.0e^{-\bar{z}/1.5})$ and with efficient spin-up modeled similar to the stellar-mass case. The resulting diagram shows the imprints of these evolutionary pathways in the mass–spin space, which cluster in regions over time and reveal mass–spin gaps (e.g.’s in the top-left, middle, and bottom right portions of the diagram) analogous to current mass gaps such as the pair instability and neutron star-black hole gaps.
@@
The flow of trajectories toward the upper right of the plot simulate a black hole analogue of the stellar main sequence which we name Cosmic Accretion, i.e., a universal growth pathway via coherent accretion over cosmic time. We note that canonical population III stars [93], if included in Fig. 3, would highly overlap the stellar-mass and intermediate-mass distributions with very uncertain natal spins and evolve toward the supermassive regime through Cosmic Accretion; whereas supermassive population III stars [94, 95] that form from gravitational collapse of atomically-cooled primordial halos would serve as a heavy seed source of supermassive black holes and occupy the mass–spin gap in the lower right portion of Fig. 3. Further populations can be added to the diagram to extend its utility for specific use cases, such as other direct collapse black hole formation possibilities or primordial black holes.
@@
Figure 4: The predictions of three models for the spin magnitude of the Wolf-Rayet star tidally synchronized with a black hole companion in the isolated formation channel. The masses $m_{2}$ and spin magnitudes $\chi_{2}$ correspond to those of the secondary black holes (i.e., formed from the initially less massive star) in binaries from the COMPAS population [96] with color corresponding to the semi-major axis $a_{\rm preSN2}$ prior to its formation in a supernova explosion. The three spin models are indicated by triangles (Qin et al. (2018)), circles (Bavera et al. (2020)) and x’s (Steinle & Kesden (2021)). The models of Qin et al. (2018) and Bavera et al. (2020) and of Steinle & Kesden (2021) are numerical and analytical extensions of the model of [97], respectively.
@@
  • In cosmological simulations tracking SMBH assembly via galaxy mergers, the mass–spin history encodes growth mode and redshift evolution [104].
  • For intermediate-mass black holes, observed at high redshift or in dwarf galaxies, the diagram can diagnose whether their mass–spin signatures follow stellar or SMBH-like growth.

### v4

@@ I Introduction
Figure 1 displays an HR diagram from a synthetic stellar population generated with the COMPAS [^1] COMPAS is a publicly available population synthesis code designed for evolution prescriptions of isolated binary-star systems. isolated-binary evolution code with default parameter settings [6]. The stellar classifications, as the HR diagram reveals, are intimately tied to the phases of stellar evolution through the stars’ lifespans. We filter this prototypical theoretical population to obtain a small randomized subset whose evolutionary stages are shown in Fig. 1. In chronological order, these intermediate-to-high mass stars begin core hydrogen fusion, i.e. the main sequence (red dashed lines) and move through a very brief Hertzsprung gap characterized by shell helium burning, which is followed by drastic radial expansion as core helium fusion ignites and constitutes the dominant component of the stellar luminosity in the giant branch (dot-dash purple lines). Then these stars lose their envelopes, either through binary mass transfer processes or stellar wind mass loss, revealing the core-helium burning phase of stellar evolution (dotted green lines) which leads imminently to gravitational collapse, possibly associated with a supernova explosion and compact object formation. A few surviving white dwarfs (gray dots) and neutron stars (blue stars) occupy their respective lower luminosity regions with the neutron stars effective temperature being much higher due to differences in their supernovae mechanisms. This highlights the utility of the HR diagram for organizing the variety of stellar types and diagnosing stellar evolution models. Extensions of the HR diagram are still being found for varied purposes [7, 8, 9, 3, 10], underscoring its impact in modern astronomy. However, black holes could only be drawn on the HR diagram if they are embedded in an accretion disk or due to non-classical possibilities such as Hawking radiation.
@@
Motivated by this paradigm, we ask: can an analogous diagram be constructed for astrophysical black holes? We propose that, similar to the HR diagram, one can visually explore how black holes can be organized and classified. While black holes are historically challenging to observe, their known macroscopic properties of mass and spin from General Relativity offer natural axes for classification. With electromagnetic (EM) techniques such as X-ray spectroscopy and radio interferometry providing constraints on stellar-mass ($\approx 5-10^{2}\rm\,M_{\odot}$), intermediate-mass ($\approx 10^{2}-10^{5}\rm\,M_{\odot}$), and supermassive ($\gtrsim 10^{6}\rm\,M_{\odot}$) black holes [11, 12, 13, 14, 15, 16] and with gravitational-wave (GW) detections now delivering catalogs of stellar-mass black hole mergers [17, 18, 19], we can consider how patterns in black hole properties may emerge.
@@
Recent years have seen substantial effort in quantifying and interpreting the mass–spin relations of BHs across the mass spectrum. For stellar‑mass black holes in X‑ray binaries, two primary techniques—continuum‑fitting and iron‑line reflection spectroscopy—have yielded a range of spin magnitudes, underscoring the complexity of natal spin and accretion histories [20, 21]. At the supermassive end, X‑ray reflection analyses show that the majority with masses below $\approx$3\text{\times}{10}^{7}$\rm\,M_{\odot}$ are rapidly spinning ($\chi\gtrsim$ 0.9), while those with higher masses exhibit a tentative trend toward lower spins [13]. The elusive intermediate‑mass black holes remain less certain, but numerous strong candidates are found in globular clusters, hyper‑luminous X‑ray sources, and dwarf-galaxy nuclei, albeit with large uncertainties [22, 23], and a few have been confidently detected via GWs, such as GW190521 [24] and GW231123 [25], calling into question the mass boundaries that typify this mass regime. These and other classes of black holes are summarized in Table 1 along with their observed mass and spin ranges which we use to construct the datasets used in Fig.’s 3, 5, 6, 7 and in Table 2 of Sections II and III.
Table 1: Summary of the three main mass regimes/classes of black holes from current electromagnetic (EM) [26, 13] and gravitational-wave (GW) [27] observations. Only stellar-mass black holes have been observed via both EM and GW messengers. The corresponding observed ranges of mass and spin measurements of each regime are indicated along with the methods for obtaining them: X-ray reflection spectroscopy (or iron-line method), spectral continuum fitting, reverberation mapping, and stellar dynamics.
BH class/name Messenger Mass range $\rm\,M_{\odot}$ Spin range Observation methods
Stellar-mass EM $\approx 5-20$ $\approx 0.1-0.98$ reflection, continuum, dynamics
GW $\approx 5-250$ $\approx 0-0.998$ unmodeled and modeled searches
Intermediate-mass EM $\approx 10^{2}-10^{5}$ ? reflection, continuum, reverberation, dynamics
Supermassive EM $\gtrsim 10^{6}$ $\approx 0.6-0.99$ reflection, continuum, reverberation, dynamics
@@
Emergent gravitational‑wave population studies suggest a mass‑dependent spin distribution in merging binaries [28], and prior studies have explored black hole mass, spin, and redshift relationships in GW data [29, 30, 31]. An observed anti-correlation in the $\chi_{\rm eff}$$q$ plane [32, 33, 34], where $\chi_{\rm eff}$ is the aligned effective spin parameter [35, 36] and $q$ the mass ratio, was subsequently confirmed in more recent analyses of gravitational-wave catalogs and helps reveal subpopulations in the observed dataset associated with different formation scenarios [37, 38, 39]. Structures can be inferred from the spin distributions using data-driven models that capture unanticipated features without inserting systematic bias [40]. Investigations into the influence of stellar rotation on black hole mass–spin outcomes has long been studied and is model dependent [41, 42], which often involves implicitly projecting onto the mass–spin plane. These works significantly advance understanding of stellar-mass black hole populations, and motivate development of a tool for interpreting black hole mass and spin measurements.
@@
To begin sketching this framework, we present a cartoon schematic of the mass–spin plane containing a feature labeled “Cosmic Accretion” that represents a clustering of black holes, analogous to the stellar main sequence, defined by extended coherent accretion episodes over cosmic time. While there is only one main sequence for stars, i.e., the dominant band of stars on the HR diagram where they spend most of their lives, we propose that black holes may have many depending on if and how they grow in mass and spin. We quantify this by sampling from physically motivated mass and spin distributions across the stellar-mass, intermediate-mass, and supermassive black hole regimes, and modeling their redshift evolution under Eddington-limited accretion to trace evolutionary tracks/trajectories across the diagram. Next, we construct a mass–spin diagram focused on the isolated-binary formation channel [43] of stellar-mass black holes using three state-of-the-art prescriptions for obtaining high black hole spin magnitude from tidal synchronization of its Wolf-Rayet progenitor’s spin, demonstrating the model-specific regions that a population of black holes can occupy in a canonical formation scenario.
@@
This paper is organized as follows. We present the theoretical motivation for the black hole mass-spin diagram and illustrative examples to show its utility in Sec. II. In Sec. III we apply machine learning classification pipelines to mass–spin datasets. In Sec. IV we summarize our analysis and findings, and in Sec. V we discuss the implications of our results with emphasis on future investigations.
Figure 2: A sketch of the mass–spin diagram for astrophysical black holes with dimensionless spin angular momentum (i.e. $\chi=c|\mathbf{S}|/Gm^{2}$) on the vertical axis. Regions are colored, as indicated in the legend, according to a selection of formation channels of black holes in single, binary, or higher systems across the mass spectrum: binary stars in isolation, dense stellar clusters, canonical population III stars, and hierarchical galaxy mergers. In reality, these formation channels will overlap and their prevalence will depend on redshift. The two green vertical patches with red hatch symbolize the commonly used boundaries between the stellar-mass, intermediate-mass, and supermassive regimes. Typical astrophysical processes that drive the mass and spin growth across the diagram are written in each regime, with small arrows indicating uncertainty in the region borders for mass and spin predictions of formation channels. The solid blue line extending from the stellar-mass regime to the upper right corner of the diagram represents a possible main-sequence analogue for black holes from sustained accretion episodes over cosmic time.
@@ II Black hole mass-spin diagrams
Black holes cannot meaningfully be placed on an HR diagram as they lack intrinsic luminosity but a mass–spin diagram provides an analogous classification tool. In this space, every black hole occupies a point whose coordinates can shift over cosmic time due to formation, accretion, merger events, or even Hawking evaporation, thereby tracing out trajectories (or tracks) that reflect astrophysical and cosmological history. The mappings between redshift and environment-dependent accretion states, and between pre- and post-merger binary component masses and spins, imply the existence of black hole main sequence analogues for dominant growth pathways of mass and spin. Understanding such mappings allows inference of black hole evolutionary typology across cosmic time.
@@
Modern EM techniques have placed increasingly precise limits on spin magnitudes, both in X-ray binaries and quasars/AGN [12, 13, 62]. Measurements of X-ray binaries can in principle constrain the spin orientation of the accreting black hole but the orbital inclination introduces burdensome systematics. Furthermore, it is an active area of research to develop new methods for EM observations to confidently identify dual- or multi-AGNs [63] and to distinguish a single supermassive black hole from a supermassive black-hole binary within a galactic nucleus [14], indicating a bright future for EM measurements of black hole spin orientations relative to a companion or other environmental aspects.
Black holes are observed confidently in two main mass regimes: stellar-mass black holes through electromagnetic methods, such as accreting binaries [12, 26] or astrometry, and through the terrestrial gravitational-wave detector network LVK; and supermassive black holes in galactic nuclei and quasars, observed via multi-wavelength spectroscopy, gas and stellar dynamics, and reverberation mapping [64, 65, 13, 66, 67, 26], and hopefully via gravitational waves with a future space-based detector such as LISA via binary or minor mergers [68]. The intermediate-mass regime is populated with many strong candidates [22, 23] but has not yet been confidently confirmed beyond the few hundred solar mass extraordinary events from LVK. While these LVK events confirm the existence of black holes with masses greater than 100 $\rm\,M_{\odot}$ [69, 25], it remains to be seen how the rest of the gap, i.e. from $\sim$1000 to $\sim 10^{5}$, is filled.
@@
Across the black-hole mass spectrum, several formation channels populate regions or trajectories in the mass–spin plane, which can be distinct or highly overlapping depending on the underlying assumptions. Stellar core collapse can produce stellar-mass black holes with low to moderate natal spins depending on the nature of angular momentum transport and possible tidal evolution during progenitor evolution [74, 75]. Binary mergers in dense environments lead to spin-up and mass growth via well-characterized remnant formulae. Accretion-driven evolution from both disk feeding and tidal disruptions can also significantly affect black hole spins depending on the Eddington fraction of the accretor and the tidal radius of the disruptor, respectively. Supermassive black holes grow through galaxy mergers and prolonged accretion episodes, resulting in mass–spin evolution that is skewed toward high spin in low-mass supermassive black holes but possibly limiting at high mass [76, 77, 78]. Also primordial black holes which, if they exist, could span a wide range of masses and spins and serve as fossils for cosmology.
Figure 3: Theoretical evolutionary tracks of accreting black holes across the mass–spin diagram. Black holes are drawn from initial mass functions and simple natal spin prescriptions, and are evolved from redshift $z=10$ to $z=0$ along the trajectories. Stellar (blue dots and lines) and supermassive (red triangles and lines) black holes undergo an approximate thin disk evolution, while intermediate-mass black holes (green squares and lines) experience chaotic accretion modeled with a Gaussian drift. The clusters that result from the flow of points toward the upper-right corner of the panel is a quantitative example of a Cosmic Accretion main sequence (i.e., the blue solid line of Fig. 2).
@@
In Figure 2 we present a schematic mass–spin diagram summarizing some main theoretical possibilities and expected trajectories. Each shaded region corresponds to a specific black hole formation channel, such as isolated stellar binaries, stellar clusters, hierarchical galaxy mergers, or Population III stars, and spans a range of spin magnitudes and black hole masses predicted by each scenario. Horizontal and vertical transitions are overlaid to indicate astrophysical processes relevant for growing the black hole masses and/or spins. The processes for intermediate-mass black holes, which may be populated by multiple overlapping formation pathways, are expected to be a mixture of processes identified from the stellar (i.e. stellar collapse, spin inheritance, tidal spin-up of the progenitor, and accretion) and supermassive (dynamical friction, binary loss-cone stellar scattering, disk-driven migration and accretion) regimes [22, 23]. The commonly used boundaries between the three mass regimes are shown for reference in green regions with red hatching. The small arrows colored by formation channel indicate uncertainties on the theoretical boundaries of the formation channel. Therefore, this diagram serves as a visual analog to the HR diagram for stars: a two-dimensional landscape in which different populations occupy distinct, physically motivated regions, and evolutionary tracks and snapshots can be mapped and classified. The solid blue line extending across the diagram signifies a conservative upper limit on the spins of black holes undergoing Eddington-limited accretion. The flow of tracks toward the upper-right corner motivate a black hole main sequence analogue we refer to as Cosmic Accretion, i.e., a cosmological period of coherent accretion growth. We note that super-Eddington accretion sources will produce tracks to higher spins for a given mass.
@@
We quantitatively illustrate this in Figure 3 which shows the redshift-dependent trajectories of black holes across the mass–spin plane under different physical growth scenarios, representing stellar-mass black holes (blue circles), intermediate-mass black holes (green squares), and supermassive black holes (red triangles). Each line, i.e. trajectory, traces the evolution of an individual black hole from redshift $z=10$ to $z=0$, with markers indicating the initial position in mass–spin space. The stellar-mass black holes are born with a power-law mass distribution ($m\sim m^{-2.3}$ [81]) and uniform spin magnitudes [82, 83], and evolve under an approximation to the classic model of [84] where we assume $z_{i}\rightarrow z_{j}$ with $i<j$ defines a redshift timestep. In this approximate model, a black hole of mass $m$ accretes an amount of mass $\Delta m(z)=m(z_{j})-m(z_{i})=m_{\rm n}(1+1.5(1-\bar{z}))$ where $\bar{z}=z/10$ and $m_{\rm n}$ is the natal mass, and undergoes corresponding spin increase via $\Delta\chi(z)=\chi_{i}+(\chi_{\rm eq}-\chi_{i})\eta\Delta m(z)/m$ where $\chi_{\rm eq}=0.998$ and $\eta=0.3$ is an ad-hoc parameter that controls the efficiency of spin-up. Intermediate-mass black holes are drawn from a log-flat mass distribution and bimodal initial spin distribution [85], and undergo modest stochastic mass growth $\Delta m(z)=m_{\rm n}(1+1.5(1-\bar{z}))(1+\mathcal{N}(0,0.003))$ accompanied with conditional spin drift (i.e., aggressive upward spin drift if $\chi<0.6$ at a given $z$), loosely modeling the combined effect of hierarchical mergers and chaotic accretion [86, 22]. We assume supermassive black holes originate from a 60%/40% mixture of light and heavy seed populations (light seeds from a Schechter-like distribution with natal spins $\sim\mathcal{N}(0.5,0.2)$ and heavy seeds from a log-uniform distribution with natal spins from $\sim\mathcal{N}(0.8,0.1)$) [87, 88, 77, 89, 90, 91, 92], and grow through coherent thin-disk accretion via $\Delta m(z)=m_{\rm n}(1+3.0e^{-\bar{z}/1.5})$ and with efficient spin-up modeled similar to the stellar-mass case. The resulting diagram shows the imprints of these evolutionary pathways in the mass–spin space, which cluster in regions over time and reveal mass–spin gaps (for example in the top-left, middle, and bottom right portions of the diagram) analogous to current mass gaps such as the pair instability and neutron star-black hole gaps.
@@
The flow of trajectories toward the upper right of the plot shows an analogue of the stellar main sequence for black holes that we illustrated in Fig. 2 and which we named the Cosmic Accretion main sequence where black holes born in gaseous environments (such as high-redshift galaxies) can spend a significant amount of time. Utilizing realistic initial redshifts and better models for redshift evolution could lead to different clustering in the diagram. We note that canonical population III stars [96], if included in Fig. 3, would highly overlap the stellar-mass and intermediate-mass distributions with uncertain natal spins and evolve toward the supermassive regime through the Cosmic Accretion main sequence; whereas supermassive population III stars [97, 98] that form from gravitational collapse of atomically-cooled primordial halos would serve as a heavy seed source of supermassive black holes and possibly occupy the mass–spin gap in the lower right portion of Fig. 3. Further populations can be added to the diagram to extend its utility for specific use cases, such as other direct collapse black hole formation possibilities or primordial black holes.
@@
Figure 4: The predictions of three models for the spin magnitude of the Wolf-Rayet star tidally synchronized with a black hole companion in the isolated formation channel. The masses $m_{2}$ and spin magnitudes $\chi_{2}$ correspond to those of the secondary black holes (i.e., formed from the initially less massive star) in binaries from the COMPAS population [99] with color corresponding to the semi-major axis $a_{\rm preSN2}$ prior to its formation in a supernova explosion. The three spin models are indicated by triangles (Qin et al. (2018)), circles (Bavera et al. (2020)) and x’s (Steinle & Kesden (2021)). The models of Qin et al. (2018) and Bavera et al. (2020) and of Steinle & Kesden (2021) are numerical and analytical adaptations of the model of [100], respectively.
@@
  • In cosmological simulations tracking supermassive assembly via galaxy mergers, the mass–spin history encodes growth mode and redshift evolution [107].
  • For intermediate-mass black holes, observed at high redshift or in dwarf galaxies, the diagram can diagnose whether their mass–spin signatures follow stellar or supermassive-like growth.